OpenStudy (anonymous):

x - 1 less than or equal to 7 2x greater than or equal to 22 what could these looks like on a number line, with a closed or open circle ?

OpenStudy (phi):

x -1 < = 7 to "simplify" add +1 to both sides x -1 + 1 < = 7 + 1 can you simplify both sides?

OpenStudy (anonymous):

7 + 1 = 8. x = 8.

OpenStudy (phi):

ok but you should keep the "relation" < = (which means less than or equal) (you could use \( \le\) ) x < = 8 x could be equal to 8, so put a filled in dot at 8 then show the numbers less than 8 (which side of 8 is that on a number line?

OpenStudy (anonymous):

to the left, yes?

OpenStudy (phi):

yes, to the left. a solid dot at 8 and an arrow pointing to the left (toward numbers smaller than 8)

OpenStudy (phi):

is this a separate problem: 2x greater than or equal to 22 2x > = 22

OpenStudy (anonymous):

okay. so x could be 11, yes?

OpenStudy (phi):

yes

OpenStudy (phi):

I would divide both sides by 2 to get x >= 11 solid dot at 11 (because x could = 11) and an arrow pointing to numbers bigger than 11 (x can be 11 or bigger)

OpenStudy (anonymous):

to the right side, because the numbers are larger, yes? thank you very much!

OpenStudy (phi):

good job

OpenStudy (phi):

the only tricky part is when you write > the "big end" is next to the bigger number and the "small, pointy end" is next to the smaller number so 2 > 1 or 1 < 2 both mean the same thing, and both are true.

OpenStudy (anonymous):

cool, i didn't know that! thank you :)